40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 18 12:52:00 2008, Alberto, Configuration, SUS, Mode Cleaner Cavity Alignment 
    Reply  Thu Dec 18 19:15:54 2008, Alberto, Yoichi, Configuration, SUS, Mode Cleaner Cavity Alignment 
Message ID: 1193     Entry time: Thu Dec 18 19:15:54 2008     In reply to: 1192
Author: Alberto, Yoichi 
Type: Configuration 
Category: SUS 
Subject: Mode Cleaner Cavity Alignment 

Quote:
This morning I found the MC locked to the 10 mode. When I locked it on the 00 mode, it was unstable and eventually it always got locked to the wrong mode.

I looked at the Drift Mon MEDM screen, which shows a reference record for position, pitch and yaw of each mirror, and I found that the MC optics were in a different status. Moving the sliders of the mirrors' actuators, I brought them back to the reference position. Then the lock got engaged and it was stable, although the MC reflection from the photodiode, with the wave front sensors (WFS) off, was about 2V. That's higher than the 0.5V the it could get when we aligned the cavity and the input periscope last time.

With the WFS on, the reflection dropped to 0.3V and, so far, the the cavity has been stably locked.


This evening the mode cleaner was again locking on a higher mode so we tweaked the mirrors' actuators by their sliders on the MEDM screen until we improved the reflection to 0.3V.

Then we went inside and, on the AS table, we centered the beam on the wave front sensors.

Now the mode cleaner is locked, the reflection is less than 0.3V and the transmission about 3V, tha is it is in ideal conditions. We'll see if it holds.
ELOG V3.1.3-